# Microfoundation Inference for Strategic Prediction

### v1

@@ Abstract
Often in prediction tasks, the predictive model itself can influence the distribution of the target variable, a phenomenon termed performative prediction. Generally, this influence stems from strategic actions taken by stakeholders with a vested interest in predictive models. A key challenge that hinders the widespread adaptation of performative prediction in machine learning is that practitioners are generally unaware of the social impacts of their predictions. To address this gap, we propose a methodology for learning the distribution map that encapsulates the long-term impacts of predictive models on the population. Specifically, we model agents’ responses as a cost-adjusted utility maximization problem and propose estimates for said cost. Our approach leverages optimal transport to align pre-model exposure (ex ante) and post-model exposure (ex post) distributions. We provide a rate of convergence for this proposed estimate and assess its quality through empirical demonstrations on a credit-scoring dataset. [^1] Code: https://github.com/felipemaiapolo/microfoundation_inference.
@@ 1.1 Related works
The performative prediction problem was introduced by Perdomo et al. (2020). Much of the methodological focus in performative prediction is on locating performative stable or performative optimal solutions. To locate a stable solution, a paradigm of repeated retraining has been developed in multiple environments Mendler-Dünner et al. (2020); Drusvyatskiy and Xiao (2023). Furthermore, a line of prior work on this often utilizes zero’th order algorithms such as finite differences Izzo et al. (2021, 2022) or regret minimization (Jagadeesan et al., 2022; Dong et al., 2018; Chen et al., 2020). Other works instead assume white-box access to a pre-specified performative map (Levanon and Rosenfeld, 2021; Somerstep et al., 2023; Shavit et al., 2020; Lin and Zrnic, 2023). Finally, others make structural assumptions on the underlying data (e.g assuming all data is drawn from a location-scale family) (Miller et al., 2021). As pointed out in Lin and Zrnic (2023) such pre-specified maps are often misspecified leading to erroneous results. We focus on correcting this by building a robust method to infer performative maps.
@@ 2.1 Importance of microfoundation inference for downstream tasks
$T_{\theta}(X)$ $\textstyle={\arg\max}_{x^{\prime}}{\theta}^{\top}x^{\prime}-\frac{1}{2}(X-x^{\prime})^{\top}M(X-x^{\prime})=X+M^{-1}\theta\,,$ (2.3)
$\textstyle\textsc{PR}(\theta)$ $={\mathbb{E}}\big{[}\{Y-\theta^{\top}T_{\theta}(X)\}^{2}\big{]}=\sigma^{2}+\|\theta-\theta^{\star}\|_{2}^{2}+\{\theta^{\top}M^{-1}\theta\}^{2}\,.$ (2.4)
@@
PR $(\hat{\theta})-\min\nolimits_{\theta}\textsc{PR}(\theta)\geq\textsc{PR}(\hat{\theta})-\textsc{PR}(0)=\{(c-1)^{2}-1\}\|\theta^{\star}\|_{2}^{2}+c^{4}\{{\theta^{\star}}^{\top}M^{-1}\theta^{\star}\}^{2}$
@@ 3.2 Learning from ex-ante and ex-post distributions
$\mathscr{L}(\mu,\varphi,P,Q_{\theta_{1}},\dots,Q_{\theta_{m}})\triangleq\textstyle W_{2}^{2}\big{(}\mu,(\nabla\varphi)_{\#}P\big{)}+\sum_{k=1}^{m}W_{2}^{2}\big{(}\mu,(\nabla\varphi-\nabla B_{\theta_{k}})_{\#}Q_{\theta_{k}}\big{)},$
@@
where =$\stackrel{{\scriptstyle\text{d}}}{{=}}$ denotes equality in distributions. It follows that for $\theta_{1}\dots\theta_{m}$, $(\nabla\varphi^{\star})_{\#}P$, $(\nabla\varphi^{\star}-\nabla B_{\theta_{1}})_{\#}Q_{\theta_{1}},$ $\dots,(\nabla\varphi^{\star}-\nabla B_{\theta_{m}})_{\#}Q_{\theta_{m}}$, are all identically distributed. In optimization (3.1) we estimate $\varphi^{\star}$ by exploiting this equality in the distribution, where we minimize the 2-Wasserstein variance. We use the (2-)Wasserstein distance in (3.1) because agent responses may not have common support, and it also allows us to leverage algorithms for fast computation of Wasserstein barycenters to solve (3.1) in a block coordinate descent algorithm. We summarize our method in Algorithm 1.
@@ 3.3 Learning from ex-post distributions
~$\tilde{\mathscr{L}}(\mu,\varphi,Q_{\theta_{1}},\dots,Q_{\theta_{m}})\triangleq\textstyle\sum_{k=1}^{m}W_{2}^{2}\big{(}\mu,(\nabla\varphi-\nabla B_{\theta_{k}})_{\#}Q_{\theta_{k}}\big{)}\,.$
@@ 4 Theoretical analysis
A crucial step in analyzing the estimated cost is to verify whether it is identifiable from the proposed algorithms. More specifically, we investigate whether the potential functions obtained from (3.1) and (3.4) induce a Bergman divergence identical to $d_{\varphi^{\star}}$ (the true cost function), when the empirical distributions $\widehat{P},\widehat{Q}_{\theta},\dots,\widehat{Q}_{\theta_{m}}$ are substituted by their respective population versions $P,Q_{\theta},\dots,Q_{\theta_{m}}$. Since a Bergman divergence remains identical under any linear adjustment to its inducing potential, we verify whether $\Phi^{\star}\triangleq\{\varphi^{\star}(z)+\alpha_{0}+\alpha_{1}^{\top}z\}$ contains
@@
Impact of estimated the cost.
@@ Appendix A Supplementary proofs
which are unique [Villani, 2009]. Using the local strong convexity of $\mathbf{L}(\Pi^{\star}_{\gamma},\gamma)$ at $\gamma^{\star}$ and local strong convexity of $\mathbf{L}(\widehat{\Pi}_{\gamma},\gamma)$ at $\widehat{\gamma}$ we have
@@
Since both the supports $\widetilde{\mathcal{X}}_{P}$ and $\widetilde{\mathcal{Y}}$ are compact, we use the bound in Manole and Niles-Weed [2024, Theorem 2] (with $\alpha=2$) to obtain:
@@ Appendix C Experiments
In this experiment, agents can manipulate their features once a logistic regression classifier, $\theta=(\alpha,\beta)$, is published. In this context, $\alpha$ is the intercept term while $\beta$ is the slope. We assume [^3] In our experiment, $\beta$ assumes negative entries. $B_{\theta}(x)=|\beta|^{\top}x=-\beta^{\top}x$ [Perdomo et al., 2020] and consider the case of parametric cost estimation with $\varphi(x)=\frac{1}{2}x^{\top}Mx$ where $M$ is a positive definite matrix to be learned. As we show in (2.3), agents’ response in this setting is $T_{\theta}(x)=X-M^{-1}\beta$. For the main experiment, we assume $M=0.1*I_{3}$ and include more results in Appendix C.2.

### v2

@@ Abstract
Often in prediction tasks, the predictive model itself can influence the distribution of the target variable, a phenomenon termed performative prediction. Generally, this influence stems from strategic actions taken by stakeholders with a vested interest in predictive models. A key challenge that hinders the widespread adaptation of performative prediction in machine learning is that practitioners are generally unaware of the social impacts of their predictions. To address this gap, we propose a methodology for learning the distribution map that encapsulates the long-term impacts of predictive models on the population. Specifically, we model agents’ responses as a cost-adjusted utility maximization problem and propose estimates for said cost. Our approach leverages optimal transport to align pre-model exposure (ex ante) and post-model exposure (ex post) distributions. We provide a rate of convergence for this proposed estimate and assess its quality through empirical demonstrations on a credit scoring dataset.
@@ 1.1 Related works
The performative prediction problem was introduced by Perdomo et al. (2020). Much of the methodological focus in performative prediction is on locating performative stable or performative optimal solutions. To locate a stable solution, a paradigm of repeated retraining has been developed in multiple environments Mendler-Dünner et al. (2020); Drusvyatskiy and Xiao (2023). Furthermore, a line of prior work on this often utilizes zero’th order algorithms such as finite differences Izzo et al. (2021, 2022) or regret minimization (Jagadeesan et al., 2022; Dong et al., 2018; Chen et al., 2020). Other works instead assume white-box access to a pre-specified performative map (Levanon and Rosenfeld, 2021; Somerstep et al., 2023; Shavit et al., 2020; Lin and Zrnic, 2023) or make structural assumptions on the underlying data (e.g assuming all data is drawn from a location-scale family) (Miller et al., 2021). The work Tsirtsis et al. (2024) casts finding a performatively optimal model (given access to a performative map) as an optimal transport problem.
As pointed out in Lin and Zrnic (2023) such pre-specified maps are often misspecified leading to erroneous results. We focus on correcting this by building a robust method to infer performative maps.
@@ 2.1 Importance of microfoundation inference for downstream tasks
$T_{\theta}(X)$ $\textstyle={\arg\max}_{x^{\prime}}{\theta}^{\top}x^{\prime}-\frac{1}{2}(X-x^{\prime})^{\top}M(X-x^{\prime})$ (2.3)
$=X+M^{-1}\theta\,,$
$\textstyle\textsc{PR}(\theta)$ $={\mathbb{E}}\big{[}\{Y-\theta^{\top}T_{\theta}(X)\}^{2}\big{]}$ (2.4)
$=\sigma^{2}+\|\theta-\theta^{\star}\|_{2}^{2}+\{\theta^{\top}M^{-1}\theta\}^{2}\,.$
@@
PR $(\hat{\theta})-\min\nolimits_{\theta}\textsc{PR}(\theta)\geq\textsc{PR}(\hat{\theta})-\textsc{PR}(0)$
=$=\{(c-1)^{2}-1\}\|\theta^{\star}\|_{2}^{2}+c^{4}\{{\theta^{\star}}^{\top}M^{-1}\theta^{\star}\}^{2}$
@@ 3.2 Learning from ex-ante and ex-post distributions
[…]
$\textstyle W_{2}^{2}\big{(}\mu,(\nabla\varphi)_{\#}P\big{)}+\sum_{k=1}^{m}W_{2}^{2}\big{(}\mu,(\nabla\varphi-\nabla B_{\theta_{k}})_{\#}Q_{\theta_{k}}\big{)},$
@@
where =$\stackrel{{\scriptstyle\text{d}}}{{=}}$ denotes equality in distributions. It follows that for $\theta_{1}\dots\theta_{m}$, $(\nabla\varphi^{\star})_{\#}P$, $(\nabla\varphi^{\star}-\nabla B_{\theta_{1}})_{\#}Q_{\theta_{1}},$ $\dots,(\nabla\varphi^{\star}-\nabla B_{\theta_{m}})_{\#}Q_{\theta_{m}}$, are all identically distributed. In optimization (3.1) we estimate $\varphi^{\star}$ by exploiting this equality in the distribution, where we minimize the 2-Wasserstein variance. We use the (2-)Wasserstein distance in (3.1) because agent responses may not have common support, and it also allows us to leverage algorithms for fast computation of Wasserstein barycenters to solve (3.1) in a block coordinate descent algorithm (cf. see Tseng (2001) for details on convergence). We summarize our method in Algorithm 1.
@@ 3.3 Learning from ex-post distributions
[…]
$=\textstyle\sum_{k=1}^{m}W_{2}^{2}\big{(}\mu,(\nabla\varphi-\nabla B_{\theta_{k}})_{\#}Q_{\theta_{k}}\big{)}\,.$
@@ 4 Theoretical analysis
A crucial step in analyzing the estimated cost is to verify whether it is identifiable from the proposed algorithms. More specifically, we investigate whether the potential functions obtained from (3.1) and (3.4) induce a Bergman divergence identical to $c_{\varphi^{\star}}$ (the true cost function), when the empirical distributions $\widehat{P},\widehat{Q}_{\theta},\dots,\widehat{Q}_{\theta_{m}}$ are substituted by their respective population versions $P,Q_{\theta},\dots,Q_{\theta_{m}}$. Since a Bergman divergence remains identical under any linear adjustment to its inducing potential, we verify whether $\Phi^{\star}\triangleq\{\varphi^{\star}(z)+\alpha_{0}+\alpha_{1}^{\top}z\}$ contains
@@
Impact of estimating the cost.
@@ 6 Summary and discussion
7 Code Availability
The code used in this work is available at https://github.com/felipemaiapolo/microfoundation_inference (Maia Polo, 2024).
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Not Applicable]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Microfoundation Inference for Strategic Prediction:
Supplementary Materials
@@ Appendix A Supplementary proofs
which are unique (Villani, 2009). Using the local strong convexity of $\mathbf{L}(\Pi^{\star}_{\gamma},\gamma)$ at $\gamma^{\star}$ and local strong convexity of $\mathbf{L}(\widehat{\Pi}_{\gamma},\gamma)$ at $\widehat{\gamma}$ we have
@@
Since both the supports $\widetilde{\mathcal{X}}_{P}$ and $\widetilde{\mathcal{Y}}$ are compact, we use the bound in Manole and Niles-Weed (2024, Theorem 2) (with $\alpha=2$) to obtain:
@@ Appendix C Experiments
In this experiment, agents can manipulate their features once a logistic regression classifier, $\theta=(\alpha,\beta)$, is published. In this context, $\alpha$ is the intercept term while $\beta$ is the slope. We assume [^2] In our experiment, $\beta$ assumes negative entries. $B_{\theta}(x)=|\beta|^{\top}x=-\beta^{\top}x$ (Perdomo et al., 2020) and consider the case of parametric cost estimation with $\varphi(x)=\frac{1}{2}x^{\top}Mx$ where $M$ is a positive definite matrix to be learned. As we show in (2.3), agents’ response in this setting is $T_{\theta}(x)=X-M^{-1}\beta$. For the main experiment, we assume $M=0.1*I_{3}$ and include more results in Appendix C.2.